Date: Attendance score: 5 pts (on-time) 2 pts Cate) Purpose: This ab wia help define ferms Pen vari and de rendenu variabee know a .ien i e be Know behavior. Close neu eren of Conditions. Seoringa 2 Pts if goals of experiment are written in studemrs words has made an effort toanswer wty are we or less big has expied objectives from word-for-word opts if left blank or only irrelevant things have been written here. Useful mathematical formulas and/or chemical reaction equations: Seoringa 1 pt for all necessary mathematiealformulas or chen reaction equations writtens pt for partiallistofmathematical formulas or chemical reaction equations o pts if irrelevant or leftblank Pre-Lab Questions 1 pt) calculate the number of moles of coz that would be produced if g of baking soda 0.25 g completely react with a stoichiometrically equivalent amount of acetic acid. CH3Coo H 2 (1 pt) vinegar is a 5% (w/) solution of acetic acid (CHnconH) in water. How many moles of acetic acid are contained in 1.00 mL of vinegar? (Hint: if you dont remember what the %(wlv) concentration unit means, review page 1 of this lab manual) 3 (M pt) Calculate the number of moles of Co2 that would be produced if 1.00 mL of vinegar completely reacts with a stoichiometrically equivalent amount of baking soda. 91
Answer #1

1) given: 0.25 g of baking soda (NaHCO3)

Molar mass of baking soda = 84 g/mol

No. of moles of baking soda = 0.25 g / 84 g/mol = 2.97 * 10-3 mol

Reaction fo baking soda with acetic acid

NaHCO3 + CH3COOH ----> NaC2H3O2  + H2O + CO2

1 mole of NaHCO3 will react with stoichometric amount of acetic acid to produce 1 mole of CO2

2.97 * 10-3 mol of NaHCO3 will produce  2.97 * 10-3 mol of CO2

Answer  2.97 * 10-3 mol of CO2

2) given vinegar is 5 % (w/v) it means 5 g of vinegar in 100 ml of solution

hence there will be 0.05 g in 1 ml of solution

molar mass of vinegar = 60 g/mol

so moles of Vinegar = 0.05 g /60 g/mol = 8.33 * 10-4 mol of vinegar

Answer; 8.33 * 10-4 mol of vinegar

3) accoriding to reaction stoichometry 1 mole of vinegar will produce 1 mole of CO2

so 8.33 * 10-4 mol of vinegar of will produce 8.33 * 10-4 mol of CO2

Answer :   8.33 * 10-4 mol of CO2